# A Fast and Accurate Approach for Covariance Matrix Construction

### v1

@@ Abstract
Reichel (2025) defined the bariance as $\operatorname{Bariance}(x)=\tfrac{1}{n(n-1)}\sum_{i<j}(x_{i}-x_{j})^{2}$, which admits an $\mathcal{O}(n)$ reformulation using scalar sums. We extend this to the covariance matrix by showing that $\operatorname{Cov}(\mathbf{X})=\tfrac{1}{n-1}\!\left(\mathbf{X}^{\top}\mathbf{X}-\tfrac{1}{n}\mathbf{s}\,\mathbf{s}^{\top}\right)$ with $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$ is algebraically identical to the pairwise-difference form yet avoids explicit centering. Computation reduces to a single $p\times p$ outer matrix product and one subtraction. Empirical Benchmarks in Python show clear runtime gains over numpy.cov and in non-BLAS tuned settings. Faster Gram routines such as RXTX [16] and for $\mathbf{X}\mathbf{X}^{\top}$ further reduce total cost.
@@ 1 Introduction
We also reference faster Gram-matrix routines such as RXTX [16] and for $\mathbf{X}\mathbf{X}^{\top}$, which can serve as a drop-in improvement or computational investment. Overall, the proposed estimator performs better in non–BLAS-tuned environments while preserving full analytical equivalence. The paper closes with a discussion of uses in standard-error computation and applications or generalized least-squares (GLS) estimation.
@@ 6 Cost model and where time is saved
The centered formulation first computes the mean vector $\bar{\mathbf{x}}$, forms the centered matrix $\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}$ (reading and writing $O(np)$ doubles), and then multiplies the result. The bariance-style formulation computes $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$, the Gram matrix $\mathbf{G}=\mathbf{X}^{\top}\mathbf{X}$, and performs one $p\times p$ outer product and one subtraction:
@@ 9 Benchmark figures
Figure 1: Runtime vs. $n$ for $p=10$. Warm-up, IQR trimming, and 95% bootstrap bands.
Figure 2: Runtime vs. $n$ for $p=50$. Same protocol.
Figure 3: Runtime vs. $p$ at fixed $n=4000$.
Figure 4: Bariance-only runtime for tall ($p=10$) and wide ($p=200$) matrices across several $n$.
Figure 5: Relative speed up ratios for $p=10$: centered/bariance and NumPy/bariance. Values >1>\!1 favor the bariance form.
Figure 6: Numerical agreement: $\max_{k,\ell}|\Sigma_{\text{Bar}}-\Sigma_{\text{Ctr}}|$ for $p=10$ across many draws. Values stay at floating-point noise.
Figure 7: Fixed-effects score-matrix covariance timing, warmed and IQR-cleaned.
10 Numerical equivalence in finite precision
For each simulated matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, compute
was recorded for a range of $(n,p)$ combinations. Across all tested sizes, the maximum difference stayed below $10^{-12}$ in IEEE 754 double precision, consistent with rounding and cancellation limits reported in [10]. These findings confirm that the algebraic equality between the bariance-style and centered forms holds to full double-precision accuracy.

### v2

# High-Performance Variance–Covariance Matrix Construction Using an Uncentered Gram Formulation

@@ Abstract
Reichel (2025) defined the bariance as a pairwise-difference measure that can be rewritten in linear time using only scalar sums. We extend this idea to the covariance matrix by showing that the standard matrix expression involving the uncentered Gram matrix and a correction term is algebraically identical to the pairwise-difference definition while avoiding explicit centering. The computation then reduces to one outer product of dimension p-by-p and a single subtraction. Benchmarks in Python show clear runtime gains, especially when BLAS optimizations are absent. Optionally faster Gram-matrix routines such as RXTX (Rybin et al., 2025) rurther reduce overall cost.
@@ 1 Introduction
We also reference faster Gram-matrix routines such as RXTX [16] and for $\mathbf{X}\mathbf{X}^{\top}$, which can serve as a drop-in improvement or computational investment. Overall, the proposed estimator performs better in non–BLAS-tuned environments while preserving full analytical equivalence. Here, “non–BLAS-tuned settings” refers primarily to computing environments in which NumPy dispatches to a generic or lightly optimized BLAS implementation. Examples include: (a) Python installations relying on the reference BLAS, (b) some Linux distributions where NumPy links to a single-thread OpenBLAS by default, and (c) cloud or container environments where optimized vendor BLAS (MKL, BLIS, OpenBLAS multi-thread) is not present. In these cases, the standard numpy.cov path incurs avoidable allocations and repeated passes over memory, whereas the Gram+outer-product formulation minimizes these costs. The paper closes with a discussion of uses in standard-error computation and applications or generalized least-squares (GLS) estimation.
Computational motivation.
The closed-form Gram–based covariance estimator is particularly beneficial in settings where (i) explicit centering becomes memory–bandwidth–bound because the construction of the centered matrix requires reading and rewriting an entire $n\times p$ array, or (ii) BLAS-level optimization is unavailable or only partially available (as is often the case in default Python/NumPy builds linked to reference BLAS or single-thread OpenBLAS). In these environments, the Gram–outer-product formulation typically performs fewer memory moves and concentrates computation in highly optimized level-3 BLAS kernels, yielding noticeable speedups even for moderately sized problems
@@ 6 Cost model and where time is saved
The centered formulation first computes the mean vector $\bar{\mathbf{x}}$, forms the centered matrix $\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}$ (reading and writing $\mathcal{O}(np)$ doubles), and then multiplies the result. The bariance-style formulation computes $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$, the Gram matrix $\mathbf{G}=\mathbf{X}^{\top}\mathbf{X}$, and performs one $p\times p$ outer product and one subtraction:
@@
Asymptotic cost comparison
Table 1 summarizes the main computational differences between the centered and Gram-based formulations. FLOP counts are reported in leading terms, ignoring lower-order additive contributions.
Method FLOPs (leading) Memory reads/writes Notes
Centered form $(X-1\bar{x}^{\top})$ $2np+2np^{2}$ Read/write $n\times p$ twice Requires materializing centered matrix
Gram + outer product $2np^{2}+p^{2}$ Read $X$ once; no $n\times p$ writes Avoids full centering; BLAS-3 dominant
RXTX (when computing $XX^{\top}$) $\approx 0.95\,(2np^{2})$ Same as Gram Beneficial when $n\gg p$ or $p\gg n$
Table 1: Asymptotic comparison of centered vs. Gram-based covariance construction.
RXTX becomes beneficial when the bottleneck is the formation of $XX^{\top}$ (e.g., large $n$ and moderate $p$), reducing the multiply–accumulate count by roughly five percent and providing measurable improvements in streaming or resampling applications.
@@ 7 Benchmark protocol
Hardware and BLAS environment.
Benchmarks were performed on an Apple M2 Pro CPU (aarch64) with 16 GB unified memory, running macOS 13.0. Python 3.11.7 with NumPy 1.26.4 was used; the BLAS backend was OpenBLAS 0.3.21 (as provided by the Anaconda distribution; single-threaded unless otherwise noted). This configuration is reported to clarify the meaning of “non–BLAS-tuned” environments, as NumPy’s performance is highly dependent on the available BLAS vendor and build options.
@@
Why numpy.cov is slower.
The numpy.cov implementation allocates a centered copy of $X$ of size $n\times p$, performs multiple passes over the data, and executes some operations at Python level before dispatching into BLAS. These additional memory moves and allocations dominate runtime in large-$n$ settings and explain why the Gram-based closed form can outperform the built-in implementation even though both rely on BLAS for the underlying matrix multiplications.
@@ 8 Numerical equivalence in finite precision
In addition to the maximum absolute deviation, we also computed the Frobenius-norm discrepancy $\|\Sigma_{\mathrm{Bar}}-\Sigma_{\mathrm{Ctr}}\|_{F}$ and the relative error $\|\Sigma_{\mathrm{Bar}}-\Sigma_{\mathrm{Ctr}}\|_{F}/\|\Sigma_{\mathrm{Ctr}}\|_{F}$. Both metrics behaved similarly, remaining at the level of floating-point rounding noise.
A brief numerical remark: when variables have extremely large magnitudes (e.g. on the order of $10^{12}$ or more), both centered and uncentered formulations may experience cancellation, though the Gram-form estimator may avoid an entire subtraction step and can be slightly more stable in practice.
@@ 9 Benchmark figures
All runtimes shown in Figures 19 are wall-clock times averaged over post-trimmed repetitions. Where comparisons are meaningful, y-axes are placed on consistent scales, and 95% bootstrap percentile intervals are reported uniformly across panels.
Figure 1: Runtime vs. $n$ for $p=10$. Times are trimmed wall-clock means with 95% bootstrap percentile bands after warm-up and IQR trimming.
Figure 2: Runtime vs. $n$ for $p=50$. Same trimmed wall-clock protocol with 95% bootstrap bands.
Figure 3: Runtime vs. $p$ at fixed $n=4000$. Times are trimmed wall-clock means with IQR outlier removal.
Figure 4: Bariance-only runtimes for tall ($p=10$) and wide ($p=200$) matrices across several $n$. Times represent trimmed wall-clock means.
Figure 5: Relative speed ratios for $p=10$: centered/bariance and NumPy/bariance using trimmed mean runtimes. Values >1>\!1 favor the bariance form.
Figure 6: Numerical agreement for $p=10$: maximum entrywise deviation $\|\Sigma_{\text{Bar}}-\Sigma_{\text{Ctr}}\|_{\infty}$. Differences remain at floating-point noise levels.
Figure 7: Numerical agreement for $p=10$: Frobenius-norm deviation $\|\Sigma_{\text{Bar}}-\Sigma_{\text{Ctr}}\|_{F}$. Values reflect expected rounding and cancellation limits.
Figure 8: Numerical agreement for $p=10$: relative Frobenius error $\|\Sigma_{\text{Bar}}-\Sigma_{\text{Ctr}}\|_{F}/\|\Sigma_{\text{Ctr}}\|_{F}$. Numerical differences remain negligible.
Figure 9: Fixed-effects score-matrix covariance timings for large $n$. Times are trimmed wall-clock means after warm-up and IQR cleaning.
@@ 10 Applications
Illustrative fixed-effects example.
Consider a single observational unit observed for $T=4$ periods with regressor matrix
$X_{i}=\begin{pmatrix}2&1\\ 3&1\\ 5&4\\ 4&2\end{pmatrix}$.
The column sums are
$\mathbf{s}_{i}=X_{i}^{\top}\mathbf{1}_{4}=\begin{pmatrix}2+3+5+4\\[2.0pt] 1+1+4+2\end{pmatrix}=\begin{pmatrix}14\\[2.0pt] 8\end{pmatrix}$.
Step 1: Gram matrix.
The within-unit Gram matrix is
$G_{i}=X_{i}^{\top}X_{i}=\begin{pmatrix}54&30\\ 30&22\end{pmatrix}$,
obtained from
$54=2^{2}+3^{2}+5^{2}+4^{2},\qquad 22=1^{2}+1^{2}+4^{2}+2^{2},\qquad 30=2\cdot 1+3\cdot 1+5\cdot 4+4\cdot 2.$
Step 2: Closed-form within-unit covariance.
The fixed-effects (demeaned) covariance matrix can be computed solely from aggregated quantities using
$\widehat{\Sigma}_{i}=\frac{1}{T-1}\!\left(G_{i}-\frac{1}{T}\mathbf{s}_{i}\mathbf{s}_{i}^{\top}\right).$
The outer product of column sums is
$\mathbf{s}_{i}\mathbf{s}_{i}^{\top}=\begin{pmatrix}196&112\\ 112&64\end{pmatrix},\qquad\frac{1}{T}\mathbf{s}_{i}\mathbf{s}_{i}^{\top}=\begin{pmatrix}49&28\\ 28&16\end{pmatrix}$.
Subtracting from the Gram matrix yields
$G_{i}-\frac{1}{T}\mathbf{s}_{i}\mathbf{s}_{i}^{\top}=\begin{pmatrix}5&2\\ 2&6\end{pmatrix}$.
Dividing by T1=3T-1=3 gives the within-unit covariance:
$\widehat{\Sigma}_{i}=\begin{pmatrix}\frac{5}{3}&\frac{2}{3}\\[4.0pt] \frac{2}{3}&2\end{pmatrix}$.
Step 3: Verification via explicit demeaning.
Let the column means be
$\bar{x}_{i}=\tfrac{1}{4}\mathbf{s}_{i}=(3.5,\;2).$
Demeaning each row of $X_{i}$ produces
$X_{i}-\mathbf{1}_{4}\bar{x}_{i}=\begin{pmatrix}-1.5&-1\\ -0.5&-1\\ 1.5&2\\ 0.5&0\end{pmatrix}$.
The sample covariance of the demeaned matrix is
$\frac{1}{T-1}(X_{i}-\bar{X}_{i})^{\top}(X_{i}-\bar{X}_{i})=\frac{1}{3}\begin{pmatrix}5&2\\ 2&6\end{pmatrix}=\widehat{\Sigma}_{i}.$
Thus, the Gram-based expression reproduces the fixed-effects covariance exactly, without requiring access to the raw period-level data. This illustrates the usefulness of the aggregated approach in settings where data are stored in compressed, privacy-protected, or distributed form.
@@
A further advantage of the Gram-formulation is that it naturally accommodates settings in which the analyst has access only to aggregated sufficient statistics (e.g. $X^{\top}X$ and $X^{\top}1$), as arises in federated learning, privacy-restricted databases, or large-scale panel systems. In such cases, the covariance can be computed without reconstructing individual observations.